OpenStudy (anonymous):

What are the solutions of the equation? z² – 6z – 27 = 0

OpenStudy (anonymous):

3, 9 3, –9 –3, 9 –3, –9

OpenStudy (anonymous):

z² - 6z - 27 = 0 Factor: (x - 9)(x + 3) = 0 Therefore: x = 9 and -3

OpenStudy (anonymous):

(z-9)(z+3) = 0 z = 9, -3